# LLaNA: Large Language and NeRF Assistant

### v1

@@ 5.2 NeRF captioning
Table 5: NeRF zero-shot classification.
@@ 5.3 NeRF single-round Q&A
In the single-round Q&A experiment, we test the ability of the assistants to provide precise answers to specific questions about the object instead of open-ended general descriptions. We prompt the models with the NeRF, or the image/cloud extracted from it, followed by one of the questions in the single-round Q&A annotations associated with the NeRF. We then collect the answer generated by the model and compare it against the ground-truth answer with the selected metrics. Results are reported in Tab. 4. Interestingly, PointLLM [77] performs better than LLaVA [44] in this task, likely because it has been specifically trained to answer detailed questions about objects represented as point clouds. Nevertheless, similarly to the detailed description results, LLaNA is the top-performing method across all metrics, again by large margins. This result suggests that the meta-encoder and the projector can extract fine-grained information from the NeRF, even if they are processing directly NeRF weights. Remarkably, the amount of information they can extract lets LLaNA answer more precisely than when images or point clouds are extracted from the NeRF. Indeed, as shown in the third row of Fig. 4 which reports a qualitative example, the only assistant able to answer correctly to a precise question about the appearance of the tyres of the car is LLaNA. Another qualitative result confirming the ability of LLaNA to provide high-quality answers to specific questions, in this case in a multi-round Q&A experiment, is reported in Fig. 5.
@@ 5.4 Zero-shot NeRF classification
Finally, we compare assistants on the task of zero-shot classification. We query the models with the sentence "What is the class of the NeRF/image/cloud? Choose among these: <Shapenet_classes>" where <Shapenet_classes> are the 10 ShapeNet classes available in our dataset. We consider the answer correct only if the ground truth class appears in the response. We report results in Tab. 5 on the ShapeNeRF–Text dataset. Notably, LLaNA achieves the best performance, demonstrating the effectiveness of the proposed architecture in realizing a NeRF assistant.
@@ 6 Limitations and future directions
Despite the promising results of our framework, it is the first study in this direction and several limitations are yet to be addressed. First, the pre-trained nf2vec encoder, having been trained exclusively on synthetic data from ShapeNet, may not generalize well to real-world objects. To address this, future work should create a NeRF–Text dataset including a more diverse set of objects, like the ones provided by Objaverse [13]. Another limitation is that nf2vec currently processes only MLPs, restricting our model to MLP-only NeRFs. However, with the rapid advancements in meta-networks, it may become very soon possible to extend LLaNA to more complex NeRF architectures, such as InstantNGP [52]. For instance, the approach by Lim et al. [42] suggests the feasibility of processing various input network architectures, although it is currently limited to small networks. Finally, our framework has been tested solely on object-centric NeRFs. Expanding its application to NeRFs representing entire scenes would be a compelling direction for future research.
@@ 7 Concluding remarks
This paper addressed the novel task of creating a language assistant for NeRF. We have tackled this problem by leveraging recent advances in MLLMs and meta-networks processing neural fields. We have shown that it is feasible and effective to process directly the weights of a NeRF to project it into the embedding space of tokens of the selected LLM. We have built and made publicly available a dataset of textual annotations of NeRFs and have shown that our approach compares favourably on its test set with respect to several foundational MLLMs used as baselines for the novel tasks of brief and detailed captioning, question answering, and zero-shot classification of NeRFs.
@@ A.1 Architecture
An instance of the employed NeRFs consists of a multi-layer perceptron (MLP) that contains three hidden layers, each with $64$ neurons. The ReLU activation function is applied between all layers except for the last one, which calculates the density and RGB values directly without any activation function. A frequency encoding [50] is applied to input 3D coordinates to improve the NeRF reconstruction quality. NeRFs do not take in input the view direction. The MLP processes an input coordinate $\mathbf{p}\in\mathbb{R}^{3}$, to produce a 4-dimensional vector containing the $RGB\sigma$.
@@ A.2 Training
Training a NeRF consists of minimizing the error between the rendered images from the NeRF and the ground truth images. Our NeRFs were trained using an $L_{1}$ loss between the predicted and ground truth RGB pixel intensities, weighting background pixels less than foreground pixels ($0.8$ foregrounds vs. $0.2$ background). Image rendering involves querying the neural network by feeding it 3D coordinates to obtain RGB color values and density estimates. By integrating these outputs along camera rays using volumetric rendering techniques [50], colors and opacities have been accumulated to produce the final rendered image. Each NeRF is trained until it reaches a good reconstruction quality, approximately for $2000$ steps.
@@ A.3 Generating images and point clouds from NeRFs
To compare with 2D and 3D MLLMs on the new tasks of NeRF captioning and NeRF Q&A, we need to render images or reconstruct point clouds from the NeRF. To render images, we employ the same volumetric rendering procedure used during the NeRF’s training. To convert NeRF into a point cloud, the marching cubes algorithm is first applied to the volumetric density field derived from the NeRF. This process generates a mesh by identifying isosurfaces within the density field. The mesh is then converted into a point cloud by considering only the mesh vertices, uniformly distributed in the 3D space. We sample RGB values from NeRF for each point coordinate to approximate point cloud colors. An example of data extracted from NeRF is depicted in Fig. 6.
@@
Generating images and point clouds requires the user to make some decisions, the effects of which on the assistant’s performance are not easy to anticipate. In the case of images, it is first of all difficult to decide the rendering viewpoint. It might happen that the object is not clearly visible in the picked viewpoint or that important elements are missing. Another decision is the resolution of the generated image, which, if too coarse, may prevent the identification of fine-grained details. The same concerns regarding the resolution also apply to point clouds. Yet, the modality encoder may not handle large resolutions or may greatly increase the processing time. Another important point is the additional computational time required to extract data from NeRF. For instance, extracting point clouds from NeRF with only $8192$ points requires approximately $620$ms. Moreover, the time for sampling the MLP and running a marching cube algorithm scales cubically with the desired spatial resolution. On the other hand, the time required to process the MLP weights is independent of the spatial resolution.
@@ B.1 Architecture
The nf2vec encoder consists of 4 linear layers with $512$, $512$, $512$, and $1024$ output neurons each. The final embedding extracted for a given MLP is a $1024$-dimensional array. Each linear layer is followed by a batch normalization and a ReLU activation function.
@@ Appendix C Details on ShapeNeRF–Text dataset
Instruction Prompts
@@
Appendix D Ground truth images and point clouds
This section presents the results of an experiment in which the baseline 2D and 3D MLLM models have been provided with ground-truth input images and point clouds extracted from the original 3D meshes in the dataset rather than from the NeRFs. This scenario estimates an upper bound for the performance of such approaches when used as NeRF assistants, by simulating perfect extraction of images or point clouds from the NeRFs. In other words, it simulates the ideal scenario in which the encoding of information inside a NeRF is lossless, a non-realisitc situation in which the baselines can achieve their best performance. Tab. 8, Tab. 9, and Tab. 10 show the results of this experiments on the tasks of brief description, detailed description, and single-round Q&A, respectively. For brevity, the best-performing 2D model, i.e., LLaVA [44] (on front views) and the best-performing 3D model, i.e., PointLLM [77], have been tested in this scenario. The results demonstrate that, even in this idealized and most favorable scenario for the baselines, LLaNA outperforms them.
@@ Appendix D Ground truth images and point clouds
LLaVA-vicuna-13b Image (Front View) 68.61 67.99 17.48 23.08 27.03

### v2

@@ 5.1 Foundation models as baselines
Figure 5: NeRF multi-round Q&A example.
5.2 NeRF captioning
We test the assistants’ ability to describe the NeRF content in the captioning experiments. We prompt them with the NeRF, or the image/cloud extracted from it, followed by the question which has been paired with its ground-truth description, as detailed in Section 4.2, e.g. "What’s the content of this NeRF/image/cloud?". We then collect the answers generated by the models and compare them with the ground-truth description according to the selected metrics.
Brief description.
We report results for the brief description tasks on ShapeNeRF–Text and the HST dataset in Tab. 2 and Tab. 2, respectively. Comparing LLaNA with the baselines described in Sec. 5.1, we appreciate how LLaNA achieves the best performance in most metrics, often by large margins against runner-ups. For instance, for the Sentence-BERT similarity on the ShapeNeRF–Text dataset, LLaNA achieves $68.63$, $7.63$ points more than LLaVA-vicuna13b, even if LLaNA uses a smaller LLM. Results on the HST dataset, which provides ground-truth descriptions validated by humans, are generally lower for all methods. Yet, LLaNA provides again the best performance according to most metrics. The difference in the quality of the brief description provided by LLaNA compared to the baselines is showcased by the qualitative result reported in the first row of Fig. 4, where the description provided by LLaNA is the most accurate.
A clear trend in both tables and qualitative results is that image-based models tend to perform better than models processing point clouds. This is likely due to the larger amount of data used during training of the modality encoder, i.e. millions of images versus hundreds of thousands of shapes, which enhances their generalization ability, as well as the capability of images to capture more details than point clouds at the input resolutions required by image-based MLLMs versus 3D MLLMs. Nonetheless, our method, which operates on NeRFs, benefits from a holistic view of the object and provides the most accurate descriptions. Remarkably, in LLaNA, all the necessary information for this language task can be extracted from a single global embedding obtained by directly processing the NeRF weights. It is also worth pointing out that, while LLaNA directly processes weights and thus is independent by design from spatial resolution, the baselines face a computational overhead growing with the desired resolution due to the necessity of extracting spatial data from NeRF (Sec. A.3). Results show that 3D-LLM performs better than the point-based models and comparably to image-based models. Comparing the results of image-based MLLMs when processing front versus back views, we can see that the vantage point has a non-negligible effect on the performance of such baselines, with SentenceBERT and SimCSE metrics diminishing by about 4 points in all baselines. In a dataset without canonical poses for objects, this would be a relevant limitation that processing NeRF weights seamlessly sidesteps. Finally, we observe that the multi-view setup of LLaVA provides similar performance to the single-view counterpart. In Appendix G, additional qualitative examples are provided.
Detailed description.
We evaluate the performance for the detailed description tasks on the proposed ShapeNeRF–Text, reporting the results in Tab. 3. For this task, the point-based model PointLLM [78] performs similarly to the image-based one, LLaVA [44]. However, we appreciate that LLaNA achieves the best performance in all metrics by large margins. For instance, for the Sentence-BERT metric, LLaNA achieves $77.43$, notably $18.35$ points more than LLaVA-vicuna-13b single-view and $17.22$ for the LLaVA-vicuna-13b multi-view setup. These large improvements indicate that, while individual or aggregated images may be sufficient for brief descriptions, they may lack all the details needed to provide a comprehensive description. Moreover, the dependency of the output quality on the selected vantage points remains strong. Contrarily, the NeRF weights contain detailed and complete information about the object, which is fundamental for more granular description tasks, with the additional advantage of not requiring tuning such hyperparameters. The ability of NeRF to capture holistic information about the object is also shown in the second row of Fig. 4, where only the direct processing of NeRF weights lets LLaNA understand that the object is a TV. PointLLM and LLaVA provide detailed but wrong descriptions, likely because of the need to extract the intermediate discrete representation as a point cloud or an image, losing information. Indeed, in both cases, it is hard even for a human observer to provide the right description from the intermediate modalities shown in the figure. More qualitative examples of this task are shown in Appendix G.
@@ 5.3 NeRF single-round Q&A
In the single-round Q&A experiment, we test the ability of the assistants to provide precise answers to specific questions about the object instead of open-ended general descriptions. We prompt the models with the NeRF, or the image/cloud extracted from it, followed by one of the questions in the single-round Q&A annotations associated with the NeRF. We then collect the answer generated by the model and compare it against the ground-truth answer with the selected metrics. Results are reported in Tab. 4. Interestingly, PointLLM [78] performs better than LLaVA [44] in this task, likely because it has been specifically trained to answer detailed questions about objects represented as point clouds. Nevertheless, similarly to the detailed description results, LLaNA is the top-performing method across all metrics, again by large margins. This result suggests that the meta-encoder and the projector can extract fine-grained information from the NeRF, even if they are processing directly NeRF weights. Remarkably, the amount of information they can extract lets LLaNA answer more precisely than when images or point clouds are extracted from the NeRF. Indeed, as shown in the third row of Fig. 4 which reports a qualitative example, the only assistant able to answer correctly to a precise question about the appearance of the tyres of the car is LLaNA. In Appendix G, additional qualitative examples of this task are provided. Finally, another qualitative result confirming the ability of LLaNA to provide high-quality answers to specific questions, in this case in a multi-round Q&A experiment, is reported in Fig. 5.
Table 6: NeRF brief captioning on ShapeNeRF-Text. Trained baselines.
Best results are in bold, runner-up is underlined.
(FV: front-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 42.86 43.22 15.56 13.74 15.27
PointLLM-7b Point cloud 55.48 57.28 21.67 25.84 24.54
GPT4Point-Opt-2.7b Point cloud 37.96 39.00 21.33 22.29 24.88
LLaNA-7b NeRF 68.63 70.54 20.64 28.33 31.76
Table 7: NeRF brief captioning on the HST dataset. Trained baselines.
Best results are in bold, runner-up is underlined.
(FV: front-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 33.79 42.66 10.28 13.22 12.19
PointLLM-7b Point cloud 44.65 44.68 8.91 12.33 12.64
GPT4Point-Opt-2.7B Point cloud 30.50 31.08 8.12 12.35 11.62
LLaNA-7b NeRF 55.62 55.56 6.56 11.81 14.52
Table 8: NeRF detailed captioning on ShapeNeRF-Text. Trained baselines.
Best results are in bold, runner-up is underlined.
(FV: front-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 44.69 42.31 10.08 23.46 12.70
PointLLM-7b Point cloud 67.30 59.56 15.39 21.42 11.37
GPT4Point-Opt-2.7b Point cloud 41.33 40.52 14.48 19.15 13.80
LLaNA-7b NeRF 77.43 79.81 41.32 36.18 32.39
Table 9: NeRF single-round Q&A on ShapeNeRF-Text. Trained baselines.
Best results are in bold, runner-up is underlined.
(FV: front-view)
Model Modality S-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 56.29 62.36 26.87 29.55 30.49
PointLLM-7b Point cloud 79.24 80.38 46.00 52.60 42.36
GPT4Point-Opt-2.7b Point cloud 22.22 28.66 8.76 13.46 14.19
LLaNA-7b NeRF 81.03 81.56 46.16 53.17 50.15
@@ 5.4 Zero-shot NeRF classification
Finally, we compare assistants on the task of zero-shot classification. We query the models with the sentence "What is the class of the NeRF/image/cloud? Choose among these: <Shapenet_classes>" where <Shapenet_classes> are the 10 ShapeNet classes available in our dataset. We consider the answer correct only if the ground truth class appears in the response. We report results in Sec. 5.1 on the ShapeNeRF–Text dataset. Using multiple views boosts the zero-shot classification performance of LLaVA, which turns out to be the best model for this task, followed by LLaNA.
5.5 Training baselines on ShapeNeRF–Text
Tabs. 7, 7, 9 and 9 report results on language tasks of several baselines trained on ShapeNeRF–Text, while Tab. 13 of the appendix, shows zero-shot NeRF classification performance of such models. We employed those baselines on ShapeNeRF–Text, for which we were able to run the official training code. Accordingly, we followed their protocol, which, for all of them, keeps the modality-specific encoder frozen and trains an adaptor and the LLM in two steps. We notice that the trained baselines exhibit different behaviors to their frozen counterparts, with LLaVA performing significantly worse and PointLLM showing clear improvements. As for GPT4Point, we observe greater variability across metrics; however, overall, it shows no significant benefit from training on ShapeNeRF–Text. LLaNA yields the best performance compared to all baselines, either frozen or trained on ShapeNeRF–Text. Finally, Appendix F shows the generalization performance on Objaverse of LLaNA and the trained baselines.
@@ 6 Limitations and future directions
Despite the promising results of our framework, it is the first study in this direction and several limitations are yet to be addressed. First, the pre-trained nf2vec encoder, having been trained exclusively on synthetic data from ShapeNet, may not generalize well to real-world objects. To address this, future work should create a NeRF–Text dataset including a more diverse set of objects, like the ones provided by Objaverse [13] and OmniObject3D [76]. Another limitation is that nf2vec currently processes only MLPs, restricting our model to MLP-only NeRFs. However, with the rapid advancements in meta-networks, it may become very soon possible to extend LLaNA to more complex NeRF architectures, such as InstantNGP [52]. For instance, the approach by Lim et al. [42] suggests the feasibility of processing various input architectures, although it is currently limited to small networks. Finally, our framework has been tested solely on object-centric NeRFs. Expanding its application to NeRFs representing entire scenes would be a compelling direction for future research.
@@ 7 Concluding remarks
This paper addressed the novel task of creating a language assistant for NeRF. We have tackled this problem by leveraging recent advances in MLLMs and meta-networks processing neural fields. We have shown that it is feasible and effective to directly process the weights of a NeRF to project it into the input embedding space of an LLM. We have built and made publicly available a dataset of textual annotations of NeRFs and have shown that our approach compares favourably with respect to several MLLMs used as baselines for the novel tasks of brief and detailed captioning, question answering, and zero-shot classification of NeRFs.
@@ A.1 Architecture
An instance of the employed NeRFs consists of a multi-layer perceptron (MLP) that contains three hidden layers, each with $64$ neurons. The ReLU activation function is applied between all layers except for the last one, which calculates the density and RGB values directly without any activation function. A frequency encoding [50] is applied to the input 3D coordinates, in order to improve the NeRF reconstruction quality. NeRFs do not take as input the view direction. The MLP processes an input coordinate $\mathbf{p}\in\mathbb{R}^{3}$, to produce a 4-dimensional vector containing $RGB\sigma$.
@@ A.2 Training
Training a NeRF consists of minimizing the error between the rendered images from the NeRF and the ground truth images. Our NeRFs were trained using an $L_{1}$ loss between the predicted and ground truth RGB pixel intensities, weighting background pixels less than foreground pixels ($0.8$ foregrounds vs. $0.2$ background). Image rendering involves querying the neural network by feeding it 3D coordinates to obtain RGB color values and density estimates. By integrating these outputs along camera rays using volumetric rendering techniques [50], colors and opacities are accumulated to produce the final rendered image. Each NeRF is trained until it reaches a good reconstruction quality, approximately for $2000$ steps.
@@ A.3 Generating images and point clouds from NeRFs
To compare with 2D and 3D MLLMs on the new tasks of NeRF captioning and NeRF Q&A, we need to render images or reconstruct point clouds from the NeRF. To render images, we employ the same volumetric rendering procedure used during the NeRF’s training. In order to extract a point cloud, the marching cubes algorithm is first applied to the volumetric density field derived from the NeRF. This process generates a mesh by identifying isosurfaces within the density field. The mesh is then converted into a point cloud by considering only the mesh vertices, uniformly distributed in the 3D space. We sample RGB values from NeRF for each point coordinate to approximate point cloud colors. An example of data extracted from NeRF is depicted in Fig. 6.
@@
Generating images and point clouds requires the user to make some decisions, the effects of which on the assistant’s performance are not easy to anticipate. When dealing with images, it is difficult to select the rendering viewpoint. It might happen that the object is not clearly visible from the chosen viewpoint or that important elements are missing. Another decision is the resolution of the generated image, which, if too coarse, may prevent the identification of fine-grained details. The same concerns regarding the resolution also apply to point clouds. Yet, the modality encoder may not handle large resolutions or may greatly increase the processing time. Another important point is the additional computational time required to extract data from NeRF. For instance, extracting point clouds from NeRF with only $8192$ points requires approximately $620$ms. Moreover, the time for sampling the MLP and running a marching cube algorithm scales cubically with the desired spatial resolution. On the other hand, the time required to process the MLP weights is independent of the spatial resolution.
@@ B.1 Architecture
The nf2vec encoder consists of 4 linear layers with $512$, $512$, $512$, and $1024$ output neurons each. The final embedding extracted for a given MLP is a $1024$-dimensional array. Each linear layer is followed by a batch normalization and a ReLU activation function. The weight matrices and biases of the input NeRF are stacked along the row dimension to form a matrix $\mathbf{M}\in\mathbb{R}^{S\times H}$, where $S=144+1+L*(H+1)+H+1=L*H+L+H+146$ and $L$ is the number of units per hidden layer, while $H$ is the dimension of the input.
@@ Appendix C Details on ShapeNeRF–Text dataset
C.1 Instruction prompts and ground-truth questions
In this section, we provide the instruction prompts used to generate the ground-truth answers of ShapeNeRF–Text and the list of questions used to build the ground-truth questions for the brief and detailed descriptions.
@@ C.1 Instruction prompts and ground-truth questions
Ground-truth questions for the brief and detailed descriptions
@@
C.2 ShapeNeRF–Text statistics
The average lengths in words of the instructions/responses are $8.51$/$22.76$ for brief descriptions, $7.82$/$77.90$ for detailed descriptions, $8.81$/$14.25$ for single-round QAs and $8.80$/$14.14$ (per round) for multi-round QAs. Fig. 9 and Fig. 10 report instruction/response length histograms and the word clouds obtained after removing generic words like “model”, “object” and “NeRF”, emphasizing frequent words in the ground-truth instructions and responses.
Figure 9: ShapeNeRF-Text statistics for ground-truth brief and detailed descriptions.
Brief Descriptions - Word clouds
Refer to caption Refer to caption
Instructions Responses
Brief Descriptions - Lengths (Words)
Refer to caption Refer to caption
Instructions Responses
Detailed Descriptions - Word clouds
Refer to caption Refer to caption
Instructions Responses
Detailed Descriptions - Lengths (Words)
Refer to caption Refer to caption
Instructions Responses
Figure 10: ShapeNeRF-Text statistics for ground-truth single-round and multi-round Q&A.
Single-round Q&A - Word clouds
Refer to caption Refer to caption
Instructions Responses
Single-round Q&A - Lengths (Words)
Refer to caption Refer to caption
Instructions Responses
Multi-round Q&A - Word clouds
Refer to caption Refer to caption
Instructions Responses
Multi-round Q&A - Lengths (Words)
Refer to caption Refer to caption
Instructions Responses
C.3 ShapeNeRF–Text quality analysis
We have carried out several experiments to assess the quality of the questions in ShapeNeRF–Text. More specifically, the purpose of this analysis was to understand how many questions referred to a detail that is visible only from a specific viewpoint of the object. First, we evaluated our dataset questions with a language-only model, LLaMA3. For each question $Q$, we asked LLaMA3:
Is a random viewpoint of the object enough to answer this question?
<$Q$>
If so, reply "YES"; if a specific viewpoint is needed, answer "NO".
By doing so, we obtained 5163 “YES” and 5847 “NO”, highlighting that most questions refer to some details which are visible only from a point of view.
Second, we ran a Vision-Language model, LLaVA-1.6-13b, on each question of the single-round Q&A dataset, on the front and back views of objects. Then, we selected only the LLaVA responses where the answer for the front or back view achieves a SimCSE score higher than 80%, i.e., likely correct answers, which selects approximately 45% of the answers. Among these correct responses, we calculated the percentage of those where the front and back answers are extremely different (i.e., a difference in SimCSE scores > 10). Remarkably, 26% of such answers are correct from one point of view but wrong from the other: these questions would have required multi-view information to be answered correctly. We report two qualitative examples in Fig. 11. In the first row, the Mercedes-Benz logo cannot be recognized from the back view. In the second row, from the rear viewpoint, the monitor seems turned off, and thus it is not possible to identify correctly the helicopter displayed on the screen. Similarly, Fig. 14 of the Appendix shows other examples of this kind of cases.
Figure 11: Front vs back results with LLaVA. The dataset contains many view-dependent questions.
Appendix D Additional baselines results and details
D.1 Language-only baseline
To assess potential spurious patterns in the question-answer relationships, we evaluate the performance of LLaMA 2, the LLM on which LLaNA relies, fine-tuned on ShapeNeRF-Text. In this training and evaluation protocol, the LLM is provided with questions belonging to the dataset and must return the correct answers without having access to the NeRF data. Therefore, the predicted answers may be generated only based on the textual patterns present in the training set. Results are shown in Tab. 12. A significant performance gap exists between LLaMA 2 and LLaNA, highlighting that our dataset consists of questions that can only be answered with access to information about 3D objects.
Table 12: Language-only baseline.
@@ D.1 Language-only baseline
Brief LLaMA2 Text-only 42.62 40.70 25.14 24.53 25.53
LLaNA-7b NeRF 68.63 70.54 20.64 28.33 31.76
Detailed LLaMA2 Text-only 49.73 47.68 15.15 23.27 14.78
LLaNA-7b NeRF 77.43 79.81 41.32 36.18 32.39
SingleQA LLaMA2 Text-only 68.37 68.46 44.07 51.15 48.00
LLaNA-7b NeRF 81.03 81.56 46.16 53.17 50.15
D.2 Zero-Shot NeRF classification of trained baselines
We report in Tab. 13 the results obtained on zero-shot NeRF classification task by the baselines trained on ShapeNeRF-Text. Results follow the same trend as the other language tasks, reported in the main paper.
Table 13: Zero-shot NeRF classification on ShapeNeRF-Text. Trained baselines.
Best results are in bold, runner-up is underlined. (FV: front-view)
Model Modality Accuracy (%)
LLaVA-vicuna-13b Image (FV) 36.49
PointLLM-7b Point cloud 49.69
GPT4Point-Opt-2.7b Point cloud 26.30
LLaNA-7b NeRF 67.14
Appendix E Ground-truth images and point clouds
This section presents the results of an experiment in which the baseline 2D and 3D MLLMs have been provided with ground-truth input images and point clouds extracted from the original 3D meshes in the dataset rather than from the NeRFs. This scenario estimates an upper bound for the performance of such approaches when used as NeRF assistants, by simulating perfect extraction of images or point clouds from the NeRFs. In other words, it simulates the ideal scenario in which the encoding of information inside a NeRF is lossless, a non-realisitc situation in which the baselines can achieve their best performance. Tab. 14, Tab. 15, and Tab. 16 show the results of this experiments on the tasks of brief description, detailed description, and single-round Q&A, respectively. For brevity, the best-performing 2D model, i.e., LLaVA [44] (on front views) and the best-performing 3D model, i.e., PointLLM [78], have been tested in this scenario. The results demonstrate that, even in this idealized and most favorable scenario for the baselines, LLaNA outperforms them.
Model Modality Sentence-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 68.61 67.99 17.48 23.08 27.03